Find the vector form of the equation of the plane that passes through the point P0 = (1, - 2,3) and has normal vector n =<3,1,-1>.


Find an equation for the plane that contains the line x = - 1 +3t, y=5 +3t, z =2 +t and is parallel to the line of intersection of the planes x-2(y - 1) +3z =-1 and y =-2x - 1 =0


1
Expert's answer
2021-07-15T11:25:15-0400

1)"x_o=1\\
y_o=-2\\
z_o=3\\
a=3\\
b=1\\
c=-1\\
a(x-x_o)+b(y-y_o)+c(z-z_o)=0\\
3(x-1)+1(y-(-2))-1(z-3)=0\\
3x+y-z+2=0\\
Let\ y=s,z=t\\
x=-\frac{2}{3}-\frac{1s}{3}+\frac{1t}{3}\\
y=0+1s+0t\\
z=0+0s+1t\\"

Vector form becomes

"<-\frac{2}{3},0,0>+s<-\frac{1}{3},1,0>+t<\frac{1}{3},0,1>\\"

2)

"x\u22122(y\u22121)+3z=\u22121\\\\\ny = \u22122x \u22121\\\\\ny=-2x-1\\\\\nx+4x+4+3z=-1\\\\\ny=-2x-1\\\\\nz=-\\dfrac{5}{3}x-\\dfrac{5}{3}\\\\\n<0,\u22121,\u2212\\frac{5}{3}>+t<1,\u22122,\u2212\\frac{5}{3}>\\\\\n\n\u200bThe\\ plane\\ contains\\ the\\ line\\\\\n\n\n\n\\dfrac{x-0}{1}=\\dfrac{y+1}{-2}=\\dfrac{z+\\dfrac{5}{3}}{-\\dfrac{5}{3}}\\\\\n\nAn\\ equation\\ for\\ the\\ plane\\ parallel\\ to\\\\ the\\ line\\ of\\ intersection\\ of\\ the\\ planes\\\\\n\n\n\nax+by+cz+d=0,ax+by+cz+d=0,\\\\\nwhere\\\\\n\n\n\na-2b-\\dfrac{5}{3}c=0\\\\\nThe\\ plane\\ contains\\ the\\ line\\\\\n x = \u22121 + 3t, y = 5 + 3t, z = 2 + t\\\\\n\nt=0: Point(-1, 5, 2)\\\\\n\n\nt=-2: Point(-7, -1, 0)\\\\\n\n\n\n-a+5b+2c+d=0\\\\\n-7a-b+d=0\\\\\n\n\nb=-7a+d\\\\\n\n-a-35a+5d+2c+d=0\\\\\n\n\nb=-7a+d\\\\\nc=18a-3d\\\\\nSubstitute\\\\\n\n\n\na+14a-2d-30a+5d=0\\\\\nd=5a\\\\\nIf\\ a=1\\\\\n\n\n\nd=5\\\\\nb=-2\\\\\nc=3\\\\\nThe\\ equation\\ for\\ the\\ plane\\ is\\\\\n\n\n\nx-2y+3z+5=0"